OpenStudy (sasogeek):

\(\large \frac{log24\sqrt{3}- \frac{3}{2}}{log{12}-1} \) simplify

OpenStudy (sasogeek):

\[\huge \frac{log24\sqrt{3}- \frac{3}{2}}{log{12}-1} \]

OpenStudy (turingtest):

do we know the base of the log?

OpenStudy (sasogeek):

10

OpenStudy (turingtest):

\[\frac{\log(24)\sqrt3-\frac32}{\log(12)-1}\]radical 3 is outside the root, right?

OpenStudy (turingtest):

outside the log I mean

OpenStudy (turingtest):

or is it inside the log?

OpenStudy (sasogeek):

it's inside the log

OpenStudy (sasogeek):

lemme type out the whole question.... i'm probably taking the wrong path

OpenStudy (turingtest):

I think I got it...

OpenStudy (sasogeek):

but when i type the whole question and where i've gotten to into the calculator, i get the same answer...

OpenStudy (turingtest):

\[\frac{\log(24\sqrt3)-\frac32}{\log(12)-1}=\frac{\log(2^3\cdot3^{3/2})-\frac32}{\log(2^2\cdot3)-1}\]is this at all like what you did?

OpenStudy (turingtest):

\[=\frac{\log(2^3)+\log(3^{3/2})-\frac32}{\log(2^2)+\log(3)-1}\]\[=\frac{3\log2+\frac32\log3-\frac32}{2\log2+\log3-1}\]\[=\frac{6\log2+3\log3-3}{4\log2-2\log3-2}\]not sure what more you want to do to it

OpenStudy (turingtest):

just right click and you can add whatever you forgot

OpenStudy (turingtest):

too late...

OpenStudy (sasogeek):

\[\huge \frac{log\sqrt{27}+log8-log\sqrt{1000}}{log{1.2}} \]

OpenStudy (turingtest):

I would try to break it all down as much as possible first, looking for powers to pull out of the logs

OpenStudy (turingtest):

\[\frac{\log(3^{3/2})+\log(2^3)-\log(10^{3/2})}{\log(\frac65)}\]now apply the log rules

OpenStudy (sasogeek):

i'll let u know what i get :) thanks

OpenStudy (turingtest):

welcome :D

OpenStudy (turingtest):

actually the answer is kinda neat, but you have to break it /all/ the way down

OpenStudy (turingtest):

\[\frac{\log(3^{3/2})+\log(2^3)-\log((5\cdot2)^{3/2})}{\log(\frac{3\cdot2}5)}\]/now/ apply the log rules

OpenStudy (sasogeek):

brb

OpenStudy (sasogeek):

the calculator says the answer is 3/2 but i'm still not sure which rules to apply to get there... i think i'm only applying some of them and i have limited knowledge on the rest so i'm probably not seeing it.... can u help me go down further the solution, or even rather give me the rules i'm supposed to use... :/

OpenStudy (turingtest):

sure, you only need three rules\[\log(ab)=\log a+\log b\]\[\log(\frac ab)=\log a-\log b\]\[\log(a^b)=b\log a\]I'll work it, you tell me when to stop...

OpenStudy (turingtest):

\[\frac{\log(3^{3/2})+\log(2^3)-\log((5\cdot2)^{3/2})}{\log(\frac{3\cdot2}5)}\]\[\frac{\frac32\log3+3\log2-\frac32\log(5\cdot2)}{\log(3\cdot2)-\log5}\]

OpenStudy (turingtest):

multiply by 2/2 and split up the multiples with the first log rule above\[\frac{3\log3+6\log2-3\log(5\cdot2)}{2\log(3\cdot2)-2\log5}\]\[\frac{3\log3+6\log2-3\log5-3\log2}{2\log3+2\log2-2\log5}\]\[\frac{3\log3+3\log2-3\log5}{2\log3+2\log2-2\log5}\]clear now?

OpenStudy (sasogeek):

why aren't I allowed to punch it into the calculator?

OpenStudy (turingtest):

you are, but calculators don't simplify for you...

OpenStudy (sasogeek):

well the idea of multiplying through never occurred to me... thanks

OpenStudy (sasogeek):

i meant multiplying through by 2/2

OpenStudy (turingtest):

sure thing In general it is nice to get rid of fractions in the numerator and denominator once I did that, out popped the answer

OpenStudy (sasogeek):

:) thanks a ton